OpenStudy (anonymous):

35/175 in simplest form

OpenStudy (unklerhaukus):

factor the numerator and denominator, then cancel the common factors

OpenStudy (anonymous):

so 35=5*7 and 175= 5*5*7

OpenStudy (inkyvoyd):

okay then you get \(\Huge \frac{5*7}{5*5*7}\)

OpenStudy (inkyvoyd):

so cancel out the common factors on the top and the bottom

OpenStudy (anonymous):

this is gonna sound stupid what are the common factors?

OpenStudy (inkyvoyd):

5 and 7

OpenStudy (unklerhaukus):

one of the fives, and the seven are common to the numerator And denominator, so they cancel away

OpenStudy (anonymous):

so 5 and 7

OpenStudy (inkyvoyd):

yes

OpenStudy (unklerhaukus):

\[\frac{5\times7}{5\times5\times7}=\frac{\cancel{5\times7}}{5\cancel{\times5\times7}}=\]